# {tocentry}[Uncaptioned image]

### v3

@@ 4.2 Scaled probability distributions
Such distributions are probably not realizable on real quantum hardware (except $\alpha=0$) but allow us to reason about the properties of alternative probability distributions while keeping the ranking of the CI weights intact (i.e., if $p_{I}>p_{J}$, the scaled distributions also satisfy $\tilde{p}_{I}>\tilde{p}_{J}$). In Figure 3, we show the effect of the scaled distribution for our N2 example. Letting $10\leq\alpha<$ 1 will increase the frequency of finding determinants with small weights, alleviating the sampling problem but making the resulting CI expansions less compact. Clearly, for $\alpha=0.5$ (light blue dots) in Figure 3, the sampling problem is essentially not present, with the DPS remaining high, even when going towards a large number of samples. The resulting CI expansion is, for a low number of determinants, not very good (note the high-error cyan points in the left panel of Figure 3). As the number of samples increases, most of the important determinants become selected, and a reasonably good energy is obtained, albeit with much less compact CI expansion, with around twice the number of determinants for the same energy accuracy as HCI or QSCI (with an unmodified probability distribution, $\alpha=1$). It should be noted that the (reduced) compactness of the wave function translates directly to the costs of the classical diagonalization step, with too-large expansions eventually becoming computationally intractable. In fact, we need so many determinants for $\alpha=0.5$ that the memory and compute requirements in the solution of the SCI Hamiltonian begin to become significant. Here, it should be mentioned that whereas QSCI requires only a single diagonalization step (assuming ideal execution such that configuration recovery is not required), HCI, as an iterative process, requires repeated diagonalization and search steps, which adds a computational overhead to HCI compared to QSCI for a given compactness. The cost of diagonalization depends on the size of the CI dimension, which is small in the initial HCI iterations, making the early diagonalizations computationally inexpensive. In Table S1, we provide a concrete example of the total diagonalization costs for HCI, QSCI, and the scaled QSCI distribution with alleviated sampling efficiency ($\alpha=0.5$) for a fixed target energy accuracy. Our results show that the total computational costs of HCI and the scaled QSCI distribution are comparable (when ignoring any costs related to obtaining and processing quantum samples for QSCI), which shows that the overhead in HCI is manageable. The overall scaling is favourable for HCI (see also Section 2 in the Supporting Information).
@@
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ34 and heavy-hex 2-LUCJ35 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations33, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S6. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. However, to the best of our knowledge, implementing this optimization in practice proves to be a significant challenge and would still suffer from the fundamental problems of QSCI reported above. In general, the structure and parameter initialization of UCJ/LUCJ ansätze are not yet well-understood. For example, in Ref. 43, the authors find that LUCJ circuits with completely random parameter initializations performed better than the more established approach of assigning parameters based on CCSD $t_{2}$-amplitudes. Thus, the poorer performance of LUCJ compared to exact ground-state sampling in the present study should not be taken as a definitive indication of its potential, and the effectiveness of such approximate ansätze could improve as theoretical understanding advances.
@@ 4.3 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan52, 53. We note that the same system has been studied in several QSCI works33, 40. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.
@@
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 52 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 33 and Figure S10 therein). In Ref. 33, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 33, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. We believe this discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation56 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF51. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible. Notably, our results contradict the claim that for [2Fe-2S], quantum sampling is more efficient than the classic HCI method33.
@@ 4.4 Generality and Sampling Considerations
So far, we have considered two specific systems and a class of scaled distributions for sampling. With these systems, we found a clear trade-off between sampling efficiency and the compactness of the resulting CI expansion. Strengthening the generality of our results, in the Supporting Information, we provide further examples of 29 additional chemical systems derived from the W4-17-MR benchmark set57 with some additional hand-picked systems (see Figures S8-S36). The conclusions drawn so far appear to be valid across a range of chemical systems, with the typical obstacle being poor sampling efficiency (similar to N2). Of interest are systems such as B2 and twisted ethene. These systems exhibit a behavior similar to the [2Fe-2S] system in the sense that there is an intermediate region where QSCI can provide more compact CI expansions than HCI initialized from the HF determinant, although this advantage is eventually lost when reaching more sizeable CI expansions, i.e., more accurate energies. A common feature of these systems is that the Hartree-Fock determinant is a poor initial guess for HCI as it has negligible weight in the ground-state CI expansion. Therefore, when HCI is initialized from the HF determinant, there is an intermediate region with the possibility of converging to excited states. However, as the threshold $\varepsilon$ is lowered, eventually, one connects to determinants important for the ground state, which leads to rapid convergence to the ground-state energy. These drawbacks can sometimes be avoided with good chemical intuition (recall [2Fe-2S]) or starting the HCI from determinants selected from a CIS calculation (twisted ethene).
@@
In our work, we have generally assumed that the sampling from QSCI should occur from the ground-state distribution, in line with the original proposal of QSCI30, 31. Arguably, this is also the implicit assumption in the larger-scale SQD studies33, which have employed optimization-free LUCJ with parameter initialization from CCSD $t_{2}$-amplitudes for their quantum experiments. However, the authors note that “quantum-classical optimization could further improve the quality of the solutions.” and present in the supporting information simulator studies with LUCJ circuits “optimized to minimize the estimator energy” and perform in-depth analysis on the ground state wave function concentration for quantum sampling. Optimization (i.e., energy minimization) of such ansätze would, in all likelihood, bring the probability distributions closer to the FCI distribution. Our work has demonstrated clear sampling inefficiencies when employing such distributions. However, there are several ways in which the sampling inefficiency could be a non-issue.
@@
First, if future quantum devices turn out to be orders of magnitude faster than current hardware, brute-force sampling from the ground-state distribution might be a viable strategy, although significant improvements to the performance of quantum hardware would be required. In some of the cases presented in the Supporting Information (e.g., NH3), reaching even a moderate accuracy requirement of $10^{-3}$ Hartree requires sampling counts of up to $10^{9}$. In Ref. 33, the authors report a sampling rate of three million executions for 45 minutes of QPU runtime, which translates to a sampling rate of around 1000 samples per second. The work in Ref. 33 is based on superconducting qubit technology, which can likely be considered “state-of-the-art” in terms of execution speed. With such a sample rate, performing $10^{9}$ quantum samples would take 10 days of continuous QPU execution for the sampling routine of a single calculation in an ideal, noise-free setting (excluding classic diagonalization costs). This can be considered prohibitively expensive, especially considering that the corresponding complete HCI calculations (including diagonalization costs) can be executed in about 15 minutes on a single CPU core. Thus, improvements to the execution speeds of future quantum hardware would have to be quite significant for the QSCI method to become cost-effective. Given that SQD is promoted as a current to near-term algorithm, the sampling efficiency is a serious problem.
@@
Second, one could consider alternative sampling strategies. For a fixed number of determinants ($L$), a very efficient sampling strategy would likely involve sampling from a uniform distribution over the $L$ largest CI coefficients, requiring approximately $O(2L\log(L))$ samples to achieve near-full coverage33. While we cannot rule out the existence of accurate and sample-efficient quantum circuits, their practical design and implementation are unknown and have yet to be demonstrated for QSCI. It constitutes a highly non-trivial problem and would most likely require a solution to the FCI problem in the first place to find the $L$ largest coefficients. In contrast, sampling from the ground state distribution (or at least approximations thereof) seems feasible and has already been demonstrated to some extent.
@@
A further problem arises when scaling towards larger systems. As shown in Figure 5, both QSCI and HCI incur exponentially increasing computational costs with increasing system size, which comes as no surprise, since any SCI method (independent of the sampling being done classically or quantum) inherently has an exponential computational scaling with system size58. We use linear chains of lithium hydride as an example. The bond- and intermolecular distances are selected to probe either weakly (near-equilibrium) or strongly correlated (stretched) regimes. The weakly correlated regime features a probability distribution that is similar to that of the previously discussed N2, while the strongly correlated regime has a flatter probability distribution similar to [2Fe-2S]. On a positive note, we find the sampling costs from QSCI show an exponential scaling with a smaller exponent than the corresponding complete HCI calculation time. Thus, with high-enough sampling rates, QSCI could have the potential to show usefulness.
@@
For the largest chain size considered, $(\mathrm{LiH})_{5}$, assuming a sampling rate of $10^{3}$, which is roughly representative of current capabilities33, we find that QSCI is about 500 times slower in the weakly correlated regime, and 19 times slower in the strongly correlated regime. Nevertheless, the slope (i.e, exponent) for QSCI at $\nu=10^{3}$ samples per second is lower than the HCI exponent, suggesting that a cross-over point should occur at some point, even when QSCI starts out slower than HCI. Extrapolation of the timings suggests that this will happen at $(\mathrm{LiH})_{7}$, but with a calculation cost of about 40 years. If one instead assumes a sampling rate of $10^{6}$ samples per second (i.e., three orders of magnitude faster than current rates), QSCI becomes on par with HCI in terms of timing in the weakly correlated regime, and five times faster in the more strongly correlated regime. Thus, without considerable improvements to the quantum sampling rates, it is difficult to consider QSCI a cost-effective method.

### v4

@@ 4.2 Scaled probability distributions
Such distributions are probably not realizable on real quantum hardware (except $\alpha=0$) but allow us to reason about the properties of alternative probability distributions while keeping the ranking of the CI weights intact (i.e., if $p_{I}>p_{J}$, the scaled distributions also satisfy $\tilde{p}_{I}>\tilde{p}_{J}$). In Figure 3, we show the effect of the scaled distribution for our N2 example. Letting $10\leq\alpha<$ 1 will increase the frequency of finding determinants with small weights, alleviating the sampling problem but making the resulting CI expansions less compact. Clearly, for $\alpha=0.5$ (light blue dots) in Figure 3, the sampling problem is essentially not present, with the DPS remaining high, even when going towards a large number of samples. The resulting CI expansion is, for a low number of determinants, not very good (note the high-error cyan points in the left panel of Figure 3). As the number of samples increases, most of the important determinants become selected, and a reasonably good energy is obtained, albeit with much less compact CI expansion, with around twice the number of determinants for the same energy accuracy as HCI or QSCI (with an unmodified probability distribution, $\alpha=1$). It should be noted that the (reduced) compactness of the wave function translates directly to the costs of the classical diagonalization step, with too-large expansions eventually becoming computationally intractable. In fact, we need so many determinants for $\alpha=0.5$ that the memory and compute requirements in the solution of the SCI Hamiltonian begin to become significant. As discussed above, whereas QSCI requires only a single diagonalization step (assuming ideal execution such that configuration recovery is not required), HCI, as an iterative process, requires repeated diagonalization and search steps, which adds a computational overhead to HCI compared to QSCI for a given compactness. In practice, one would likely still advise an approach for QSCI in which increasingly large diagonalisations are performed until some degree of convergence is obtained. The cost of diagonalization depends on the size of the CI dimension, which is small in the initial HCI iterations, making the early diagonalizations computationally inexpensive. In Table S1, we provide a concrete example of the total diagonalization costs for HCI, QSCI, and the scaled QSCI distribution with alleviated sampling efficiency ($\alpha=0.5$) for a fixed target energy accuracy. Our results show that the total computational cost of HCI is comparable to the diagonalization cost alone of the scaled QSCI distribution ($\alpha=0.5$), indicating that the overhead from repeated diagonalization in HCI is manageable. We also estimated the computational cost related to quantum sampling by including a conversion factor from quantum samples to seconds in the timings (i.e., a quantum sampling rate). Even with relatively optimistic estimates for the sampling rate of $10^{6}$ quantum samples per second, we find that the computational cost of HCI is lower than both the scaled ($\alpha=0.5$) and unscaled versions of QSCI.
Thus, while alternative input distributions might alleviate the sampling problem, the overall efficiency of the QSCI method would not necessarily improve if they result in less compact CI expansions.
@@
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ35 and heavy-hex 2-LUCJ36 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations34, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S6. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. Such optimization has been investigated to some extent in Ref. 34 using quantum simulators. However, this optimization is likely to bring the probability distribution closer to the ground-state distribution, which, as shown above, can still lead to inefficient sampling. To the best of our knowledge, implementing this optimization on quantum hardware in practice proves to be a significant challenge and would still suffer from the fundamental problems of QSCI reported above. In general, the structure and parameter initialization of UCJ/LUCJ ansätze are not yet well-understood. For example, in Ref. 44, the authors find that LUCJ circuits with completely random parameter initializations performed better than the more established approach of assigning parameters based on CCSD $t_{2}$-amplitudes. Thus, the poorer performance of LUCJ compared to exact ground-state sampling in the present study should not be taken as a definitive indication of its potential, and the effectiveness of such approximate ansätze could improve as theoretical understanding advances.
@@ 4.3 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan53, 54. We note that the same system has been studied in several QSCI works34, 41. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.
@@
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 53 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 34 and Figure S10 therein). In Ref. 34, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 34, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. Although the sampling component is not entirely identical, the (classically) variationally optimized LUCJ circuits employed in Ref. 34 likely approximate the ground-state distribution we consider reasonably well. Thus, we believe the discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation57 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF52. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible. Notably, our results challenge the claim that for [2Fe-2S], quantum sampling is more efficient than the classic HCI method34.
@@ 4.4 Generality and Sampling Considerations
So far, we have considered two specific systems and a class of scaled distributions for sampling. With these systems, we found a clear trade-off between sampling efficiency and the compactness of the resulting CI expansion. Strengthening the generality of our results, in the Supporting Information, we provide further examples of 29 additional chemical systems derived from the W4-17-MR benchmark set58 with some additional hand-picked systems (see Figures S8-S36). The conclusions drawn so far appear to be valid across a range of chemical systems, with the typical obstacle being poor sampling efficiency (similar to N2). Of interest are systems such as B2 and twisted ethene. These systems exhibit a behavior similar to the [2Fe-2S] system in the sense that there is an intermediate region where QSCI can provide more compact CI expansions than HCI initialized from the HF determinant, although this advantage is eventually lost when reaching more sizeable CI expansions, i.e., more accurate energies. A common feature of these systems is that the Hartree-Fock determinant is a poor initial guess for HCI as it has negligible weight in the ground-state CI expansion. Therefore, when HCI is initialized from the HF determinant, there is an intermediate region with the possibility of converging to excited states. However, as the threshold $\varepsilon$ is lowered, eventually, one connects to determinants important for the ground state, which leads to rapid convergence to the ground-state energy. These drawbacks can sometimes be avoided with good chemical intuition (recall [2Fe-2S]) or starting the HCI from determinants selected from a CIS calculation (twisted ethene).
@@
In our work, we have generally assumed that the sampling from QSCI should occur from the ground-state distribution, in line with the original proposal of QSCI31, 32. Arguably, this is also the implicit assumption in the larger-scale SQD studies34, which have employed optimization-free LUCJ with parameter initialization from CCSD $t_{2}$-amplitudes for their quantum experiments. However, the authors note that “quantum-classical optimization could further improve the quality of the solutions.” and present in the supporting information simulator studies with LUCJ circuits “optimized to minimize the estimator energy” and perform in-depth analysis on the ground state wave function concentration for quantum sampling. Optimization (i.e., energy minimization) of such ansätze would, in all likelihood, bring the probability distributions closer to the FCI distribution. Our work has demonstrated clear sampling inefficiencies when employing such distributions. However, there are several ways in which the sampling inefficiency could be a non-issue.
@@
First, if future quantum devices turn out to be orders of magnitude faster than current hardware, brute-force sampling from the ground-state distribution might be a viable strategy, although significant improvements to the performance of quantum hardware would be required. In some of the cases presented in the Supporting Information (e.g., NH3), reaching even a moderate accuracy requirement of $10^{-3}$ Hartree requires sampling counts of up to $10^{9}$. In Ref. 34, the authors report a sampling rate of three million executions for 45 minutes of QPU runtime, which translates to a sampling rate of around 1000 samples per second. The work in Ref. 34 is based on superconducting qubit technology, which can likely be considered “state-of-the-art” in terms of execution speed. With such a sample rate, performing $10^{9}$ quantum samples would take 10 days of continuous QPU execution for the sampling routine of a single calculation in an ideal, noise-free setting (excluding classic diagonalization costs). This may be considered prohibitively expensive, especially considering that the corresponding complete HCI calculations (including diagonalization costs) can be executed in about 15 minutes on a single CPU core. Nevertheless, the performance of future quantum devices remains uncertain. In Figure 5, we plot the accuracy (error relative to the CASCI reference) against the total calculation time for HCI and QSCI on the N2 and [2Fe-2S] systems. For QSCI, we show results assuming three different sample-to-seconds conversion rates, $\nu$. When assuming near-term sampling rates ($\nu=10^{3}$, magenta lines), QSCI is not competitive with HCI for either system. The nitrogen molecule is an example where the quality of the selected determinants with QSCI is roughly equal to that of HCI (see Figure 2), but sampling efficiency in QSCI is poor. Thus, while the diagonalization costs are roughly comparable, with some advantage to QSCI, since only a single round of diagonalization is required, the quantum sampling component can be costly. With faster sampling rates of $\nu=10^{6}$ (red lines), QSCI performs on par with HCI on the N2 molecule up to an accuracy of about 5 milli-Hartree, beyond which the costs from quantum sampling make QSCI more expensive overall. Assuming even faster sampling rates of $\nu=10^{9}$ (brown lines), QSCI would outperform HCI, with total calculation time reduced by roughly a factor of three for the accuracy of $10^{-4}$ Hartree. For the [2Fe-2S] system, sampling efficiency with QSCI is reasonably high, but the compactness of the resulting CI expansions is lower than the corresponding HCI expansions (see Figure 4). Thus, we find that for QSCI, the diagonalization costs become dominant. As a result, HCI remains more efficient from an accuracy-to-computation time standpoint throughout. In summary, improvements to the execution speeds of future quantum hardware would have to be quite significant for the QSCI method to become cost-effective, and for some ground-state probability distributions, any advantage seems out of reach. Given that SQD is promoted as a current to near-term algorithm, the sampling efficiency is a critical problem.
Figure 5: Total calculation time required to reach a given energy accuracy relative to the CASCI solution for HCI and QSCI on the N2 (left) and [2Fe-2S] systems (right). The number of quantum samples (QSCI) and the threshold $\varepsilon$ are varied to adjust the energy accuracy. For HCI, iteration is stopped when the search step adds fewer than 1% of the current number of determinants. The cumulative time (determinant search, Hamiltonian construction, and diagonalization) across all iterations is reported. For QSCI, the timings include a single diagonalization in addition to quantum sampling time, assuming the three indicated sampling rates $\nu$. For the iron-sulfur cluster, the cost of sampling becomes negligible with the higher sampling rates, causing the red $\nu=10^{6}$ and brown $\nu=10^{9}$ lines to overlap.
@@
Second, one could consider alternative sampling strategies. For a fixed number of determinants ($L$), a very efficient sampling strategy would likely involve sampling from a uniform distribution over the $L$ largest CI coefficients, requiring approximately $O(2L\log(L))$ samples to achieve near-full coverage34. While we cannot rule out the existence of accurate and sample-efficient quantum circuits, their practical design and implementation are unknown and have yet to be demonstrated for QSCI. It constitutes a highly non-trivial problem and would most likely require a solution to the FCI problem in the first place to find the $L$ largest coefficients. In contrast, sampling from the ground state distribution (or at least approximations thereof) seems feasible and has already been demonstrated to some extent.
@@
A further problem arises when scaling towards larger systems. As shown in Figure 6, both QSCI and HCI incur exponentially increasing computational costs with increasing system size, which comes as no surprise, since any SCI method (independent of the sampling being done classically or quantum) inherently has an exponential computational scaling with system size59. We use linear chains of lithium hydride as an example. The bond- and intermolecular distances are selected to probe either weakly (near-equilibrium) or strongly correlated (stretched) regimes. The weakly correlated regime features a probability distribution that is similar to that of the previously discussed N2, while the strongly correlated regime has a flatter probability distribution similar to [2Fe-2S]. On a positive note, we find the sampling costs from QSCI show an exponential scaling with a smaller exponent than the corresponding complete HCI calculation time. Thus, with high-enough sampling rates, QSCI could have the potential to show usefulness.
@@
For the largest chain size considered, $(\mathrm{LiH})_{5}$, assuming a sampling rate of $10^{3}$, which is roughly representative of current capabilities34, we find that QSCI is about 500 times slower in the weakly correlated regime, and 19 times slower in the strongly correlated regime. Nevertheless, the slope (i.e, exponent) for QSCI at $\nu=10^{3}$ samples per second is lower than the HCI exponent, suggesting that a cross-over point should occur at some point, even when QSCI starts out slower than HCI. Extrapolation of the timings suggests that this will happen at $(\mathrm{LiH})_{7}$, but with a calculation cost of about 40 years. If one instead assumes a sampling rate of $10^{6}$ samples per second (i.e., three orders of magnitude faster than current rates), QSCI becomes on par with HCI in terms of timing in the weakly correlated regime, and five times faster in the more strongly correlated regime. Thus, without considerable improvements to the quantum sampling rates, it is difficult to consider QSCI a cost-effective method. Moreover, because QSCI is assessed under ideal conditions here, relying on hardware advances would ultimately require it to compete with fault-tolerant algorithms that avoid the exponential scaling.